RhondaSommer (rhondasommer):

The length of the hypotenuse of a 45°-45°-90° triangle is 12 units. Find the exact value of the length of the leg. 6 units 6√2 units 6√3 units 12 √2 units

RhondaSommer (rhondasommer):

isnt it 6√3

RhondaSommer (rhondasommer):

:/ im so confused

563blackghost (563blackghost):

Since this is a 45-45-90 degree triangle that would mean that the two sides are equal so we can do this equation.... \[\frac{ 10 }{ \sqrt{2} }\]

563blackghost (563blackghost):

meant 12...

OpenStudy (anonymous):

Why not follow Pythagoras \[2x^2=12^2\] with the knowledge of equal legs.

563blackghost (563blackghost):

@RhondaSommer

RhondaSommer (rhondasommer):

but how do i get any of those choices?

563blackghost (563blackghost):

To change \(\LARGE{\frac{ 12 }{ \sqrt{2} }}\) into radical form you would think of what times 2 equals 12 well it would be 6....so \(\Large{\sqrt{2 \times 6}}\)so we would take the 6 on the outside and leave 2...so our answer would be...\(\Huge\color{red}{6\sqrt{2}}\)

RhondaSommer (rhondasommer):

ohhhh

563blackghost (563blackghost):

so you got it ^^

RhondaSommer (rhondasommer):

yep!!

563blackghost (563blackghost):

great ^^